# MIST: Towards Multi-dimensional Implicit Bias and Stereotype Evaluation of LLMs via Theory of Mind

### v1

@@ 1 Introduction
As large language models (LLMs) demonstrate increasingly sophisticated reasoning capabilities, the question of whether they possess a form of Theory of Mind (ToM) Premack and Woodruff (1978) has emerged as a central topic. ToM, also known as mentalizing, is the ability to infer the mental and emotional states of other beings. Since this capacity underpins meaningful communication and empathy, investigating its potential emergence in LLMs is a critical research endeavor with profound implications for developing safer and more cooperative AI systems Nguyen and others (2025).
A broader perspective on ToM includes the ability to model complex social structures involving individuals and groups Baker et al. (2017). As LLMs internalize knowledge from large-scale corpora, they inevitably learn the statistical distributions that reflect societal stereotypes. In this context, we can define stereotypes as learned but often flawed generalizations about a group. Bias is the subsequent failure that occurs when the model misapplies these group-level stereotypes to make judgments about an individual. This represents a systematic failure of ToM, as the model generates erroneous beliefs about a person’s intentions or competencies.
@@
2 Related Work
2.1 Theory of Mind in LLMs
The recent advancements in the reasoning and problem-solving capabilities of LLMs Wei et al. (2022), Wang et al. (2022), Huang and Chang (2023) have provoked significant scientific debate surrounding their potential for an emergent ToM. This capacity, defined as the ability to attribute and reason about the beliefs, intentions, and knowledge of others, has long been considered a hallmark of human social intelligence. Consequently, evaluating the extent to which LLMs can replicate ToM Kosinski (2023) has become a pivotal research objective, carrying profound implications for the future development of human-centered AI Zhao et al. (2023), Li et al. (2024), Liu et al. (2025b), Wang et al. (2024), Cheng et al. (2025), Liu et al. (2025a).
However, scholarly assessments of these capabilities have yielded divergent conclusions. Critical analyses posit that high performance may stem from methodological flaws in benchmark design or a reliance on superficial statistical patterns rather than genuine reasoning Wang et al. (2025), Sadhu et al. (2024). In contrast, other research demonstrates that ToM skills can be made more robust, showing that targeted training enables generalization to novel and complex tasks Lu et al. (2025). As LLMs internalize knowledge from vast corpora, they also learn flawed societal stereotypes. A critical failure of mentalizing occurs when a model misapplies these learned group-level generalizations to an individual, thereby forming distorted and erroneous beliefs about their intentions or competencies. Therefore, current research should address both the functional robustness of ToM in specific tasks and these broader systemic failures.
2.2 Stereotype Content Model
@@ 3.1 Tasks Definition
3.1.1 Word Association Bias Test
@@
3.1.2 Affctive Attribution Test
@@ 3.2 Evaluation Metrics
3.2.1 Word Association Bias Test
@@
3.2.2 Affective Attribution Test
We focus on the model’s affective attribution tendencies along two target-specific directions:
When the target group identifier belongs to (advantaged group) and the model classifies it as comedy (positive valence), it is counted as a favorable attribution.
When the target group identifier belongs to (disadvantaged group) and the model classifies it as tragedy (negative valence), it is counted as an unfavorable attribution.
@@ 3.2.2 Affective Attribution Test
3.3 Data Construction
3.3.1 Lexicon Resources Selection
In the race domain, following prior studies Acerbi and Stubbersfield (2023), Bai et al. (2025) , we designated American as the advantaged group, while the disadvantaged groups were selected from four representative regions: African, Asian, Spanish-speaking, and Arab. In the gender domain, consistent with previous work Bai et al. (2025), females were treated as the advantaged group, while males and transgender individuals were considered disadvantaged groups. The health domain was further divided into four subdomains, where the advantaged groups included non-disabled, slim, young, and mentally healthy individuals, and the corresponding disadvantaged groups were disabled, overweight, aged, and individuals with mental illness. The specific group identifiers111https://www.behindthename.com/top/ used for each advantaged and disadvantaged category are summarized as follows in Table 1.
Next, we define positive attribute words () and negative attribute words () based on the 3 dimensions of the SCM: Competence, Sociability, and Morality. The initial set of attribute words was partially derived from prior studies Bai et al. (2025). Subsequently, we further invite five scholars with expertise in psychology to refine, supplement, and evaluate the attribute word lists. The final selection of attribute words is presented in Table 2.
@@ 3.3.1 Lexicon Resources Selection
In addition to group identifiers and attribute words, we also constructed a set of neutral object words () to serve as fillers in the data templates. These neutral words were manually curated to represent inanimate and content-neutral entities that are not directly associated with any social group or stereotype. The words are presented in Table 3.
| Category | Words |
| Furniture | Table, Chair, Shelf |
| Vessel | Bottle, Plate, Cup, Box, Bag, Container |
| Tool | Pen, Key, Map, Coin, Wire, Pipe, Tool |
| Structure | Bridge, Window, Door, Frame, Fence |
| Nature | Road, Cloud, Stone, Hill, Path |
| Object | Book, Sheet, Lamp, Clock |
3.3.2 Data Template Design
For each task WABT and AAT, we design 3 sets of data templates to minimize potential unintended biases that may be introduced by specific prompt formulations. The data templates for each task are presented below.
@@ 3.3.2 Data Template Design
3.3.3 Data Generation
Specifically, for the WABT task, we construct 10 paired combinations of and (e.g., African vs. American, Asian vs. American, etc.). For each combination, we randomly sample one pair of and group identifiers, and subsequently sample 5 and 5 attribute words from the lexicons corresponding to the three stereotype content dimensions. This sampling procedure is repeated 50 times for each combination. The sampled items are then combined with 3 data templates, resulting in a total of 4,500 instances. For the AAT task, we randomly sample 500 combinations of group identifiers and neutral nouns , and combine them with 3 data templates, resulting in a total of 1,500 instances. Figure 2 presents the distribution of the generated data.
4 Experiments and Results
4.1 Evaluated Models
4.2 Evaluation results on WABT
We input 4,500 data into 8 LLMs and obtain their respective responses. For each data, we record the number of valid responses returned by the models. For each valid response, we further compute the frequency counts of 4 specific combinations: , , , and . Based on these counts, we calculate the bias score along the 3 stereotype dimensions—Competence, Sociability, and Morality, following our predefined computational formulas.
@@ 4.2 Evaluation results on WABT
Figure 3 provides a visual illustration of the bias score distributions across various social groups and stereotype dimensions for 4 LLMs with more than 1,000 valid responses.
Models Dimension Mean Std Claude-3.7-sonnet Competence 1470 -0.006 0.933 -0.263 0.792 Sociability 1482 0.415 0.830 19.273 .001 Morality 1465 -0.012 0.966 -0.493 0.622 DeepSeek-R1 Competence 178 -0.246 0.891 -3.672 .001 Sociability 183 0.320 0.787 5.492 .001 Morality 129 0.085 0.915 1.056 0.293 DeepSeek-V3 Competence 42 0.054 0.985 0.354 0.725 Sociability 17 0.180 0.945 0.763 0.456 Morality 23 0.130 0.991 0.617 0.544 Gemini-2.5-pro Competence 1351 -0.057 0.923 -2.252 0.025 Sociability 1398 0.367 0.792 17.329 .001 Morality 1352 -0.012 0.931 -0.489 0.625 GPT-4o Competence 382 -0.067 0.923 -1.423 0.156 Sociability 400 0.275 0.885 6.203 .001 Morality 419 0.176 0.956 3.770 .001 GPT-4-turbo Competence 1256 -0.063 0.953 -2.345 0.019 Sociability 1205 0.341 0.835 14.150 .001 Morality 1160 0.072 0.976 2.512 0.012 LlaMa-2-70B-Chat Competence 99 -0.039 0.698 -0.558 0.578 Sociability 113 0.256 0.657 4.124 .001 Morality 78 0.128 0.798 1.408 0.163 LlaMa-3-70B-Instruct Competence 1156 0.098 0.901 3.681 .001 Sociability 1130 0.246 0.854 9.682 .001 Morality 1095 0.065 0.920 2.337 0.020
4.3 Evaluation Results on AAT
We input 1,500 data into 8 LLMs and obtain their respective responses. For each data, we record the number of valid responses returned by the models. For each valid response, we further analyze the emotional framing chosen by the model—specifically, whether the response aligns more closely with a comedic or tragic interpretation.
Moreover, some models demonstrate particularly polarized patterns. For instance, DeepSeek-R1 and DeepSeek-V3 show overwhelmingly tragic framings across almost all groups, while LLaMa-2-70B-Chat produces a predominance of neutral responses, especially for marginalized identities.
Models Comedy Tragedy Neutrality Comedy Tragedy Neutrality Claude-3.7-sonnet 0.777 0.048 0.175 0.780 0.047 0.173 DeepSeek-R1 0.265 0.735 0.000 0.212 0.788 0.000 DeepSeek-V3 0.267 0.725 0.008 0.279 0.708 0.013 Gemini-2.5-pro 0.340 0.628 0.032 0.352 0.624 0.024 GPT-4o 0.267 0.725 0.008 0.279 0.708 0.013 GPT-4-turbo 0.532 0.465 0.003 0.452 0.547 0.001 LLaMa-2-70B-Chat 0.368 0.025 0.607 0.372 0.027 0.601 LLaMa-3-70B-Instruct 0.567 0.388 0.045 0.574 0.388 0.038
5 In-Depth Analysis of MIST
5.1 Consistently Observed: Pervasive Positive Bias in Sociability
Figure 3 provides further visual confirmation of this pattern. For models such as GPT-4-turbo, Gemini-2.5-pro, and LLaMa-3-70B-Instruct, the green lines representing the Sociability dimension extend outward across a broad spectrum of groups, including “Asian”, “African”, “Arab”, “Male”, and “Disabled”. This cross-model and cross-group consistency suggests a systematic inclination in the model behavior to portray entities as less sociable or friendly.
In the AAT task design, the models are required to make a binary attribution decision between “Comedy” and “Tragedy”. However, in practice, several models spontaneously generate a proportion of outputs labeled as “Neutrality”, which are not pre-specified in the response options. This phenomenon indicates that some models exhibit attributional avoidance or uncertainty under certain social contexts.
As shown in Table 5, models produce near-zero proportions of “Neutrality” such as Claude-3.7-sonnet, LLaMa-3-70B-Instruct, and LLaMa-2-70B-Chat. Notably, LLaMa-2-70B-Chat yield the highest rate of neutral attributions, with 60.7% for groups and 60.1% for groups, making it the most concentrated model in terms of neutral responses. In contrast, both Claude-3.7-sonnet and LLaMa-3-70B-Instruct maintain relatively lower neutral rates (ranging from 4% to 17% across both and groups), but still demonstrated certain stable neutral tendencies under specific subgroup.
This neutral output phenomenon may reflect 2 potential mechanisms: (1) when models face certain sensitive social group identifiers, internal stereotype conflicts may lead to indecisive attribution behaviors, manifesting as ambiguous or uncertain attributional avoidance; (2) alternatively, some models may have been influenced by safety-oriented alignment optimization during training, causing them to proactively avoid emotionally sensitive outputs and instead favor neutralized responses as part of a “safety regulation avoidance mechanism”.
5.4 Asymmetry: Divergent Patterns of Implicit Bias
This finding indicates that although all models exhibit group-level attribution bias at the global level, the underlying mechanisms through which such biases are expressed are not entirely consistent across models. Instead, models demonstrate divergent patterns in both the direction and magnitude of attribution. Some models predominantly exhibit “positive amplification for advantaged groups,” whereas others display “negative amplification for disadvantaged groups.”
6 Conclusion

### v2

# MIST: Towards Multi-dimensional Implicit BiaS Evaluation of LLMs for Theory of Mind

@@ Abstract
MIST: Towards Multi-dimensional Implicit BiaS
Evaluation of LLMs for Theory of Mind
@@ 1 Introduction
As Large Language Models (LLMs) display more sophisticated reasoning, whether they possess a form of Theory of Mind (ToM) has become a central question Premack and Woodruff (1978). ToM is the capacity to attribute and infer others’ intentions, beliefs, and mental states, and it requires inferences grounded in individualized, context-sensitive evidence. Under contemporary large-corpus pretraining, LLMs internalize group-level statistical regularities Baker et al. (2017). When these regularities are used to simplify social perception, they become imperfect generalizations in the form of stereotypes. When these regularities are used to simplify social perception, they become imperfect generalizations in the form of stereotypes. If such group representations are inappropriately applied to individual-level inference, substituting group traits for person-specific evidence, bias arises and leads to systemic failures in ToM as shown in Figure 1. Therefore, identifying and quantifying bias is a prerequisite for evaluating and improving the ToM capacity of LLMs.
@@ 2.1 Tasks Definition
Word Association Bias Test.
@@
Affective Attribution Test.
@@ 2.2 Evaluation Metrics
Word Association Bias Test.
@@
Affective Attribution Test.
We focus on the model’s affective attribution tendencies along two target-specific directions: (1) When the target group identifier belongs to (advantaged group) and the model classifies it as comedy (positive valence), it is counted as a favorable attribution. (2) When the target group identifier belongs to (disadvantaged group) and the model classifies it as tragedy (negative valence), it is counted as an unfavorable attribution.
@@
2.3 Data Construction
Lexicon Resources Selection.
In the race domain, following prior studies Acerbi and Stubbersfield (2023), Bai et al. (2025) , we designated American as the advantaged group, while the disadvantaged groups were selected from four representative regions: African, Asian, Spanish-speaking, and Arab. In the gender domain, consistent with previous work Bai et al. (2025), females were treated as the advantaged group, while males and transgender individuals were considered disadvantaged groups. The health domain was further divided into four subdomains, where the advantaged groups included non-disabled, slim, young, and mentally healthy individuals, and the corresponding disadvantaged groups were disabled, overweight, aged, and individuals with mental illness.
Next, we define positive attribute words () and negative attribute words () based on the 3 dimensions of the SCM: Competence, Sociability, and Morality. The initial set of attribute words was partially derived from prior studies Bai et al. (2025). Subsequently, we further invite five scholars with expertise in psychology to refine, supplement, and evaluate the attribute word lists.
In addition to group identifiers and attribute words, we also constructed a set of neutral object words () to serve as fillers in the data templates. These neutral words were manually curated to represent inanimate and content-neutral entities that are not directly associated with any social group or stereotype. Please refer to the Appendix B.1 for more details on lexicon resource selection.
Data Template Design.
For each task WABT and AAT, we design 3 sets of data templates to minimize potential unintended biases that may be introduced by specific prompt formulations. The detailed templates refer to the Appendix B.2.
Data Generation.
Specifically, for the WABT task, we construct 10 paired combinations of and (e.g., African vs. American, Asian vs. American, etc.). For each combination, we randomly sample one pair of and group identifiers, and subsequently sample 5 and 5 attribute words from the lexicons corresponding to the three stereotype content dimensions. This sampling procedure is repeated 50 times for each combination. The sampled items are then combined with 3 data templates, resulting in a total of 4,500 instances. For the AAT task, we randomly sample 500 combinations of group identifiers and neutral nouns , and combine them with 3 data templates, resulting in a total of 1,500 instances. The statistics details are shown in Appendix B.3. Moreover, some instances are shown in Appendix B.4.
3 Experiments and Results
3.1 Evaluated Models
3.2 Evaluation results on WABT
We input 4,500 data into 8 LLMs and obtain their respective responses. For each data, we record the number of valid responses returned by the models. For each valid response, we further compute the frequency counts of 4 specific combinations: , , , and . Based on these counts, we calculate the bias score along the 3 stereotype dimensions—Competence, Sociability, and Morality, following our predefined computational formulas. Model response examples are shown in Appendix C.1.
Figure 3 provides a visual illustration of the bias score distributions across various social groups and stereotype dimensions for 8 LLMs.
Models Dim. Mean Std Claude-3.7-sonnet Com. 1470 -0.006 0.933 -0.263 0.792 Soc. 1482 0.415 0.830 19.273 .001 Mor. 1465 -0.012 0.966 -0.493 0.622 DeepSeek-R1 Com. 178 -0.246 0.891 -3.672 .001 Soc. 183 0.320 0.787 5.492 .001 Mor. 129 0.085 0.915 1.056 0.293 DeepSeek-V3 Com. 42 0.054 0.985 0.354 0.725 Soc. 17 0.180 0.945 0.763 0.456 Mor. 23 0.130 0.991 0.617 0.544 Gemini-2.5-pro Com. 1351 -0.057 0.923 -2.252 0.025 Soc. 1398 0.367 0.792 17.329 .001 Mor. 1352 -0.012 0.931 -0.489 0.625 GPT-4o Com. 382 -0.067 0.923 -1.423 0.156 Soc. 400 0.275 0.885 6.203 .001 Mor. 419 0.176 0.956 3.770 .001 GPT-4-turbo Com. 1256 -0.063 0.953 -2.345 0.019 Soc. 1205 0.341 0.835 14.150 .001 Mor. 1160 0.072 0.976 2.512 0.012 LlaMa-2-70B-Chat Com. 99 -0.039 0.698 -0.558 0.578 Soc. 113 0.256 0.657 4.124 .001 Mor. 78 0.128 0.798 1.408 0.163 LlaMa-3-70B-Instruct Com. 1156 0.098 0.901 3.681 .001 Soc. 1130 0.246 0.854 9.682 .001 Mor. 1095 0.065 0.920 2.337 0.020
3.3 Evaluation Results on AAT
We input 1,500 data into 8 LLMs and obtain their respective responses. For each data, we record the number of valid responses returned by the models. For each valid response, we further analyze the emotional framing chosen by the model—specifically, whether the response aligns more closely with a comedic or tragic interpretation. Some response examples are shown in Appendix C.2.
Moreover, some models demonstrate particularly polarized patterns. For instance, DeepSeek-R1 and DeepSeek-V3 show overwhelmingly tragic framings across almost all groups, while LLaMa-2-70B-Chat produces a predominance of neutral responses, especially for marginalized identities. More extended experiment results are shown in Appendix D.
Models C. T. N. C. T. N. Claude-3.7-sonnet 0.777 0.048 0.175 0.780 0.047 0.173 DeepSeek-R1 0.265 0.735 0.000 0.212 0.788 0.000 DeepSeek-V3 0.267 0.725 0.008 0.279 0.708 0.013 Gemini-2.5-pro 0.340 0.628 0.032 0.352 0.624 0.024 GPT-4o 0.267 0.725 0.008 0.279 0.708 0.013 GPT-4-turbo 0.532 0.465 0.003 0.452 0.547 0.001 LLaMa-2-70B-Chat 0.368 0.025 0.607 0.372 0.027 0.601 LLaMa-3-70B-Instruct 0.567 0.388 0.045 0.574 0.388 0.038
4 In-Depth Analysis of MIST
4.1 Consistently Observed: Pervasive Positive Bias in Sociability
Figure 3 provides further visual confirmation of this pattern. For models such as GPT-4-turbo, Gemini-2.5-pro, and LLaMa-3-70B-Instruct, the lines representing the Sociability dimension extend outward across a broad spectrum of groups, including “Asian”, “African”, “Arab”, “Male”, and “Disabled”. This cross-model and cross-group consistency suggests a systematic inclination in the model behavior to portray entities as less sociable or friendly.
In the AAT task design, the models are required to make a binary attribution decision between "Comedy" and "Tragedy". However, in practice, several models spontaneously generate a proportion of outputs labeled as "Neutrality", which are not pre-specified in the response options. This phenomenon indicates that some models exhibit attributional avoidance or uncertainty under certain social contexts.
As shown in Table 2, models produce near-zero proportions of "Neutrality" such as Claude-3.7-sonnet, LLaMa-3-70B-Instruct, and LLaMa-2-70B-Chat. Notably, LLaMa-2-70B-Chat yield the highest rate of neutral attributions, with 60.7% for groups and 60.1% for groups, making it the most concentrated model in terms of neutral responses. In contrast, both Claude-3.7-sonnet and LLaMa-3-70B-Instruct maintain relatively lower neutral rates (ranging from 4% to 17% across both and groups), but still demonstrated certain stable neutral tendencies under specific subgroup.
This neutral output phenomenon may reflect 2 potential mechanisms: (1) when models face certain sensitive social group identifiers, internal stereotype conflicts may lead to indecisive attribution behaviors, manifesting as ambiguous or uncertain attributional avoidance; (2) alternatively, some models may have been influenced by safety-oriented alignment optimization during training, causing them to proactively avoid emotionally sensitive outputs and instead favor neutralized responses as part of a "safety regulation avoidance mechanism".
4.4 Asymmetry: Divergent Patterns of Implicit Bias
This finding indicates that although all models exhibit group-level attribution bias at the global level, the underlying mechanisms through which such biases are expressed are not entirely consistent across models. Instead, models demonstrate divergent patterns in both the direction and magnitude of attribution. Some models predominantly exhibit "positive amplification for advantaged groups,” whereas others display "negative amplification for disadvantaged groups.”
5 Related Work
Theory of Mind in LLMs.
Recent advancements in LLM reasoning Wei et al. (2022); Wang et al. (2022); Huang and Chang (2023) have spurred debate on their potential for an emergent Theory of Mind (ToM). Evaluating this capacity is now a pivotal goal for human-centered AI Kosinski (2023); Zhao et al. (2023); Li et al. (2024); Liu et al. (2025b); Wang et al. (2024); Cheng et al. (2025); Liu et al. (2025a). However, scholarly assessments diverge. Some critics argue that high performance stems from flawed benchmarks or superficial statistical patterns rather than genuine reasoning Wang et al. (2025); Sadhu et al. (2024), while other research shows ToM skills can be robustly trained to generalize Lu et al. (2025). A critical failure arises when LLMs misapply learned societal stereotypes to individuals. Therefore, research must address both the functional robustness of ToM and these broader systemic failures.
Stereotype Content Model.
6 Conclusion
7 Limitations
This study has some limitations. Our evaluation uses a synthetic dataset with specific word lists, which may not capture the full complexity of how bias manifests in real-world language. The framework is also focused on English and a limited set of social domains such as race, gender, and health. Therefore, our findings may not generalize to other languages or cultural contexts where stereotypes differ. Lastly, while our indirect tasks effectively probe for certain biases, they might not uncover all subtle forms of stereotyping that could appear in more complex, open-ended interactions with the models.
Appendix A Future Work
In the future, we plan to expand our framework to cover more social groups and different languages to see how bias varies across cultures. We also intend to create more complex evaluation tasks that mimic real-world scenarios to better understand how these models behave in practice. Further research will focus on identifying the sources of these biases within the training data and model architecture.
Appendix B Details of Dataset
B.1 Lexicon Selection
The specific group identifiers111https://www.behindthename.com/top/ used for each advantaged and disadvantaged category are summarized as follows in Table 4. The final selection of attribute words is presented in Table 5. And the neutral object words are presented in Table 3.
| Category | Words |
| Furniture | Table, Chair, Shelf |
| Vessel | Bottle, Plate, Cup, Box, Bag, Container |
| Tool | Pen, Key, Map, Coin, Wire, Pipe, Tool |
| Structure | Bridge, Window, Door, Frame, Fence |
| Nature | Road, Cloud, Stone, Hill, Path |
| Object | Book, Sheet, Lamp, Clock |
@@ B.1 Lexicon Selection
B.2 Data Construction Templates
Word Association Bias Test.
@@ B.2 Data Construction Templates
B.3 Data Statistics
For the WABT task we have 4,500 instances, and for the AAT task we have 1,500 instances. The detailed distribution of dataset across Race, Gender and Health domains is shown in Figure 6.
| Number | Percentage | |
| Total | 6000 | - |
| -WABT | 4500 | 75% |
| -AAT | 1500 | 25% |
| Domain | - | - |
| -Race | 2400 | 40% |
| -Gender | 1200 | 20% |
| -Health | 2400 | 40% |
B.4 Data Examples
In this section, we provide several concrete data examples from WABT and AAT, as shown in Figure 7. These examples cover different bias dimensions, including race, gender, and health. Each box presents representative input instances.
Appendix C Model Responses
C.1 Word Association Bias Test
The example C.1 is a valid paired example following the requirement. This example C.1 illustrates an instance of an invalid pairing. As observed, the model does not directly respond with a clear pairing strategy; instead, it adopts a seemingly random approach to avoid providing a definitive answer.
C.2 Affective Attribution Test
The example C.2 is a valid instance of AAT. The example C.2 is an invalid instance of AAT. The model refuse the classification request and responded with a vague and ambiguous reply.
Appendix D Extended Experiment Results
@@
In this section, we illustrate more experiment results. Figure 8 provides a visual illustration of the average bias score distributions across various social groups and stereotype dimensions for 8 LLMs, where blue denotes the Competence, green denotes Sociability, and red denotes Morality.
The results show that most models exhibit bias scores close to +1 for groups such as Disability, Overweight, and Mental illness. Since positive bias indicates that favorable descriptors are disproportionately assigned to advantaged groups, these findings suggest that disadvantaged groups are systematically deprived of positive semantic associations. Although the magnitude of bias varies across model families, the overall trend is consistent: GPT-4-turbo and the LLaMa series display relatively stable distributions close to neutrality but still skewed toward advantaged groups; GPT-4o shows larger fluctuations; the DeepSeek series produces more dispersed results and sometimes extreme values; Claude-3.7-sonnet and Gemini-2.5-pro fall in between, with Claude exhibiting particularly strong bias against Mental illness. Overall, these patterns indicate that large language models tend to reinforce existing social hierarchies across sensitive dimensions such as competence, sociability, and morality, allocating more positive descriptors to advantaged groups.
Figure 9 presents the distribution of emotional framings across different social groups for each LLM, where green denotes the proportion of Comedy, red denotes Tragedy, and blue denotes Neutrality.
The y-axis represents the percentage of each emotional category, with the total summing to 100% for each group. The results show that most models produce a certain proportion of neutral responses, whereas the DeepSeek series (R1 and V3) generate almost no neutral outputs, with results heavily concentrated in either "comedy" or "tragedy". This suggests that DeepSeek relies less on avoidance mechanisms in its alignment strategy and instead tends to provide more definitive judgments, even if such polarized expressions may reinforce group-related stereotypes. These results indicate implicit biases in LLM behaviors.